OpenStudy (anonymous):

Show that the point (pi/3, 1/2) is on the graph of y=cosx

OpenStudy (anonymous):

\[\large ( \frac{ \pi }{ 3 }, \frac{ 1 }{ 2 })\]

OpenStudy (anonymous):

Do i just substitute the coordinate and check if LHS=RHS ?

OpenStudy (loser66):

definitely yes